OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

the slope of the equation is 2...it says parallel line so slope of parallel is also 2 now use y−y1=m(x−x1) substitute the values y - 3 = 2(x-5) y - 3 = 2x - 10 2x - y - 10 + 3 = 0 can you simplify that?

OpenStudy (cwrw238):

the slope of the line y = mc + c is m co,pare this with y = 2x + 2 to find its slope

OpenStudy (cwrw238):

* compare

OpenStudy (cwrw238):

do you see why the slope is 2?

OpenStudy (anonymous):

thank you :)